OpenStudy (anonymous):

Okay. My friend got 7.5 and i got 16.5. what one is right? and if I did get it wrong, whatd i do wrong? 3-2m/5=6. solve for m.

OpenStudy (ghazi):

(3-2m)=6*5=30.........now you can calculate

OpenStudy (ghazi):

well on the other hand 3-6=2m/5........therefore 2m= -15 and m= -7.5

OpenStudy (anonymous):

you can turn it around to be 2m-3 and then add 3 to 30 and then divide 33 by 2....but what one is right?

OpenStudy (ghazi):

well it depends ...on your question...i did what i saw ...if you will consider it as (3-2m)/5=6 then you'll have another answer and if you'll consider 3-(2m/5)=6 you'll have another one

OpenStudy (anonymous):

-2m/5 = 3 -2m = 15 m = -7.5

OpenStudy (anonymous):

THIS IS SO CONFUSING. D:

OpenStudy (anonymous):

it depends on how you write the problem

OpenStudy (ghazi):

you've to go with either of them

OpenStudy (anonymous):

 this is what it looks like.